A wire loop with current I is placed in a uniform
Answer #1

1) The net force is zero, but the net torque is not (FALSE): The net force and the net torque are both zero.

\:

2) There are no forces exerted on the current loop (FALSE): There are forces exerted on the current loop, which are given by the equation,

\:

\vec{F}_{m}=I\Delta \vec{L}\times \vec{B}

\:

3) The net force is zero {\color{Blue} (TRUE)} : The net force is

\:

\vec{F}_{net}=\vec{F}_{right}+\vec{F}_{left}+\vec{F}_{top}+\vec{F}_{bottom}

\:

since,

\:

\vec{F}_{right}=-\vec{F}_{left}

\vec{F}_{top}=-\vec{F}_{bottom}

\:

we have,

\:

\vec{F}_{net}=\mathbf{0}

\:

4) The loop will rotate (FALSE): The net torque in the loop must be not zero to rotate, but the net torque is zero.

\:

\vec{\tau }_{net}=\vec{\tau }_{left}+\vec{\tau }_{right}+\vec{\tau }_{top}+\vec{\tau }_{bottom}

\vec{\tau }_{net}=rF _{left}(\sin 180^{\circ})+rF _{right}(\sin 180^{\circ})+0(F_{bottom})+0(F_{top})=\mathbf{0}